### v1

# Tokenization Is More Than Compression

Abstract
Tokenization is a foundational step in Natural Language Processing (NLP) tasks, bridging raw text and language models. Existing tokenization approaches like Byte-Pair Encoding (BPE) originate from the field of data compression, and it has been suggested that the effectiveness of BPE stems from its ability to condense text into a relatively small number of tokens. We test the hypothesis that fewer tokens lead to better downstream performance by introducing PathPiece, a new tokenizer that segments a document’s text into the minimum number of tokens for a given vocabulary. Through extensive experimentation we find this hypothesis not to be the case, casting doubt on the understanding of the reasons for effective tokenization. To examine which other factors play a role, we evaluate design decisions across all three phases of tokenization: pre-tokenization, vocabulary construction, and segmentation, offering new insights into the design of effective tokenizers. Specifically, we illustrate the importance of pre-tokenization and the benefits of using BPE to initialize vocabulary construction. We train 64 language models with varying tokenization, ranging in size from 350M to 2.4B parameters, all of which are made publicly available.
Tokenization Is More Than Compression
Craig W. Schmidt  Varshini Reddy  Haoran Zhang†,‡  Alec Alameddine Omri Uzan§  Yuval Pinter§  Chris Tanner†,¶ Kensho Technologies Harvard University §Department of Computer Science MIT Cambridge, MA Cambridge, MA Ben-Gurion University of the Negev Cambridge, MA Beer Sheva, Israel {craig.schmidt,varshini.reddy,alec.alameddine,chris.tanner}@kensho.com haoran_zhang@g.harvard.edu{omriuz@post,uvp@cs}.bgu.ac.il
1 Introduction
Tokenization is an essential step in NLP that translates human-readable sentences into a sequence of distinct tokens that can be subsequently used by statistical models Grefenstette (1999). Recently, a growing number of studies have researched the effects of tokenization, both in an intrinsic manner and as it affects downstream model performance Singh et al. (2019); Bostrom and Durrett (2020); Hofmann et al. (2021, 2022); Limisiewicz et al. (2023). To rigorously inspect the impact of tokenization, we consider tokenization as three distinct, sequential stages:
  1. 1.
    Pre-tokenization: an optional set of initial rules that restricts or enforces the creation of certain tokens (e.g., splitting a corpus on whitespace, thus preventing any tokens from containing whitespace).
  2. 2.
    Vocabulary Construction: the core algorithm that, given a text corpus $\mathcal{C}$ and desired vocabulary size $m$, constructs a vocabulary of tokens $t_{k}\in\mathcal{V}$, such that $|\mathcal{V}|=m$, while adhering to the pre-tokenization rules.
  3. 3.
    Segmentation: given a vocabulary $\mathcal{V}$ and a document $d$, segmentation determines how to split $d$ into a series of $K_{d}$ tokens $t_{1},\dots,t_{k},\dots,t_{K_{d}}$, with all $t_{k}\in\mathcal{V}$, such that the concatenation of the tokens strictly equals $d$. Given a corpus of documents $d\in\mathcal{C}$, we will define the corpus token count (CTC) as the total number of tokens used in each segmentation, $CTC=\sum_{d\in\mathcal{C}}K_{d}$.
    As an example, segmentation might decide to split the text intractable into “int ract able”, “in trac table”, “in tractable”, or “int r act able”.
[…]
The widely used Byte-Pair Encoding (BPE) tokenizer Sennrich et al. (2016) originated in the field of data compression Gage (1994). Gallé (2019) argues that it is effective because it compresses text to a short sequence of tokens. To investigate the hypothesis that having fewer tokens leads to better downstream performance, we design a novel tokenizer, PathPiece, that, for a given document $d$ and vocabulary $\mathcal{V}$, finds a segmentation with the minimum possible $K_{d}$. The PathPiece vocabulary construction routine is a top-down procedure that heuristically minimizes CTC on a training corpus. PathPiece is ideal for studying the effect of CTC on downstream performance, as we can vary decisions at each tokenization stage.
We extend these experiments to the most commonly used tokenizers, focusing on how downstream task performance is impacted by the major stages of tokenization and vocabulary sizes. Toward this aim, we conducted experiments by training 64 language models (LMs): 54 LMs with 350M parameters; 6 LMs with 1.3B parameters; and 4 LMs with 2.4B parameters. We provide open-source, public access to PathPiece, our token vocabularies, and all our 64 trained LMs. [^1] https://github.com/kensho-technologies
2 Related Work
2.1 Pre-tokenization Methods
Pre-tokenization is a process of breaking text into chunks, which are then tokenized independently. A token is not allowed to cross these pre-tokenization boundaries. BPE, WordPiece, and Unigram all require new chunks to begin whenever a space is encountered. If a space appears in a chunk, it must be the first character; hence, we will call this “FirstSpace”. Thus “␣New” is allowed but “New␣York” is not. Gow-Smith et al. (2022) suggest treating spaces as individual tokens, which we will call “Space” pre-tokenization, while Jacobs and Pinter (2022) suggest leaving spaces at the end of tokens. Llama Touvron et al. (2023) popularized the idea of having each digit always be an individual token, which we call “Digit” pre-tokenization.
2.2 Vocabulary Construction
We focus on byte-level, lossless subword tokenization. Subword tokenization algorithms split text into word and subword units based on their frequency and co-occurrence patterns from their “training” data, effectively capturing morphological and semantic nuances in the tokenization process Mikolov et al. (2011). Mielke et al. (2021) gives a survey of subword tokenization.
We analyze BPE, WordPiece, and Unigram as baseline subword tokenizers, using the implementations from HuggingFace [^2] https://github.com/huggingface/tokenizers with ByteLevel pre-tokenization enabled. We additionally study SaGe, a context-sensitive subword tokenizer, using the recently released 2.0 version [^3] https://github.com/MeLeLBGU/SaGe.
Byte-Pair Encoding
Sennrich et al. (2016) introduced Byte-Pair Encoding (BPE), a bottom-up method where the vocabulary construction starts with single bytes as tokens. It then merges the most commonly occurring pair of adjacent tokens in a training corpus into a single new token in the vocabulary. This process repeats until the desired vocabulary size is reached. Issues with BPE and analyses of its properties are discussed in Bostrom and Durrett (2020); Klein and Tsarfaty (2020); Gutierrez-Vasques et al. (2021); Yehezkel and Pinter (2023); Saleva and Lignos (2023); Liang et al. (2023).
WordPiece
WordPiece is similar to BPE, except that it uses Pointwise Mutual Information (PMI) Bouma (2009) as the criteria to identify candidates to merge, rather than a count Wu et al. (2016); Schuster and Nakajima (2012). PMI prioritizes merging pairs that occur together more frequently than expected, relative to the individual token frequencies.
Unigram Language Model
Unigram works in a top-down manner, starting from a large initial vocabulary and progressively pruning groups of tokens that induce the minimum likelihood decrease of the corpus Kudo (2018). This selects tokens to maximize the likelihood of the corpus, according to a simple unigram language model.
SaGe
Yehezkel and Pinter (2023) proposed SaGe, a subword tokenization algorithm incorporating contextual information into an ablation loss via a skipgram objective. SaGe also operates top-down, pruning from an initial vocabulary to a desired size.
2.3 Segmentation Methods
Given a tokenizer and a vocabulary of tokens, segmentation converts text into a series of tokens. We included all 256 single-byte tokens in the vocabulary of all our experiments, ensuring any text can be segmented without out-of-vocabulary issues.
Certain segmentation methods are tightly coupled to the vocabulary construction step, such as merge rules for BPE or the maximum likelihood approach for Unigram. Others, such as the WordPiece approach of greedily taking the longest prefix token in the vocabulary at each point, can be applied to any vocabulary. Alternative segmentation schemes include Dynamic Programming BPE He et al. (2020), BPE-Dropout Provilkov et al. (2020), and FLOTA Hofmann et al. (2022).
3 PathPiece
Several efforts over the last few years (Gallé, 2019; Zouhar et al., 2023, inter alia) have suggested that the empirical advantage of BPE as a tokenizer in many NLP applications, despite its unawareness of language structure, can be traced to its superior compression abilities, providing models with overall shorter sequences during learning and inference. Inspired by this claim we introduce PathPiece, a lossless subword tokenizer that, given a vocabulary $\mathcal{V}$ and document $d$, produces a segmentation minimizing the total number of tokens needed to split $d$. We additionally provide a vocabulary construction procedure that, using this segmentation, attempts to find a $\mathcal{V}$ minimizing the corpus token count (CTC). [^4] An extended description is given in Appendix A. PathPiece provides an ideal testing laboratory for the compression hypothesis by virtue of its maximally efficient segmentation.
3.1 Segmentation
PathPiece requires that all single-byte tokens are included in vocabulary $\mathcal{V}$ to run correctly. PathPiece works by finding a shortest path through a directed acyclic graph (DAG), where each byte $i$ of training data forms a node in the graph, and two nodes $j$ and $i$, contain a directed edge if the byte segment $[j,i]$ is a token in $\mathcal{V}$.
We describe PathPiece segmentation in Algorithm 1, where $L$ is a limit on the maximum width of a token in bytes, which we set to 16. It has a complexity of $O(nL)$, which follows directly from the two nested for-loops. For each byte $i$ in $d$, it computes the shortest path length $pl[i]$ in tokens up to and including byte $i$, and the width $wid[i]$ of a token with that shortest path length. In choosing $wid[i]$, ties between multiple tokens with the same shortest path length $pl[i]$ can be broken randomly, or the one with the longest $wid[i]$ can be chosen, as shown here. [^5] Random tie-breaking, which can be viewed as a form of subword regularization, is presented in Appendix A. Some motivation for selecting the longest token is due to the success of FLOTA Hofmann et al. (2022). Then, a backward pass constructs the shortest possible segmentation from the $wid[i]$ values computed in the forward pass.
1:procedure PathPiece($d,\mathcal{V},L$)
2:     $n\leftarrow\operatorname{len}(d)$triangleright document length
3:     $pl[1:n]\leftarrow\infty$triangleright shortest path length
4:     $wid[1:n]\leftarrow 0$triangleright shortest path tok width
5:     for $e\leftarrow 1,n$ do $\triangleright$ token end
6:         for $w\leftarrow 1,L$ do $\triangleright$ token width
7:              $s\leftarrow e-w+1$triangleright token start
8:              if $s\geq 1$ then $\triangleright$ $s$ in range
9:                  if $d[s:e]\in\mathcal{V}$ then
10:                       if $s=1$ then $\triangleright$ 1 tok path
11:                           $pl[e]\leftarrow 1$
12:                           $wid[e]\leftarrow w$
13:                       else
14:                           $nl\leftarrow pl[s-1]+1$
15:                           if $nl\leq pl[e]$ then
16:                                $pl[e]\leftarrow nl$
17:                                $wid[e]\leftarrow w$
18:     $T\leftarrow[\,]$triangleright output token list
19:     $e\leftarrow n$triangleright start at end of $d$
20:     while $e\geq 1$ do
21:         $s\leftarrow e-wid[e]+1$triangleright token start
22:         $T.\operatorname{append}(d[s:e])$triangleright append token
23:         $e\leftarrow e-wid[e]$triangleright back up a token
24:     return $\operatorname{reversed}(T)$triangleright reverse order
Algorithm 1 PathPiece segmentation.
3.2 Vocabulary Construction
PathPiece’s vocabulary is built in a top-down manner, attempting to minimize the corpus token count (CTC), by starting from a large initial vocabulary $\mathcal{V}_{0}$ and iteratively omitting batches of tokens. The $\mathcal{V}_{0}$ may be initialized from the most frequently occurring byte $n$-grams in the corpus, or from a large vocabulary trained by BPE or Unigram. We enforce that all single-byte tokens remain in the vocabulary and that all tokens are $L$ bytes or shorter.
For a PathPiece segmentation $t_{1},\dots,t_{K_{d}}$ of a document $d$ in the training corpus $\mathcal{C}$, we would like to know the increase in the overall length of the segmentation $K_{d}$ after omitting each token $t$ from our vocabulary and then recomputing the segmentation. Tokens with a low overall increase are good candidates to remove from the vocabulary.
To avoid the very expensive $O(nL|\mathcal{V}|)$ computation of each segmentation from scratch, we make a simplifying assumption that allows us to compute these increases more efficiently: we omit a specific token $t_{k}$, for $k\in[1,\dots,K_{d}]$ in the segmentation of a particular document $d$, and compute the minimum increase $MI_{kd}\geq 0$ in the total tokens $K_{d}$ from not having that token $t_{k}$ in the segmentation of $d$. We then aggregate these token count increases $MI_{kd}$ for each token $t\in\mathcal{V}$. We can compute the $MI_{kd}$ without actually re-segmenting any documents, by reusing the shortest path information computed by Algorithm 1 during segmentation.
Any segmentation not containing $t_{k}$ must either contain a token boundary somewhere inside of $t_{k}$ breaking it in two, or it must contain a token that entirely contains $t_{k}$ as a superset. We enumerate all occurrences for these two cases, and we find the minimum increase $MI_{kd}$ among them. Let $t_{k}$ start at index $s$ and end at index $e$, inclusive. Path length $pl[j]$ represents the number of tokens required for the shortest path up to and including byte $j$. We also run Algorithm 1 backwards on $d$, computing a similar vector of backwards path lengths $bpl[j]$, representing the number of tokens on a path from the end of the data up to and including byte $j$. The minimum length of a segmentation with a token boundary after byte $j$ is thus:
$K_{j}^{b}=pl[j]+bpl[j+1].$ (1)
We have added an extra constraint on the shortest path, that there is a break at $j$, so clearly $K_{j}^{b}\geq K_{d}$. The minimum increase for the case of having a token boundary within $t_{k}$ is thus:
$MI_{kd}^{b}=\min_{j=s,\dots,e-1}{K_{j}^{b}-K_{d}}.$ (2)
The minimum increase from omitting $t_{k}$ could also be from a segmentation containing a strict superset of $t_{k}$. Let this superset token be $t_{k}^{\prime}$, with start $s^{\prime}$ and end $e^{\prime}$ inclusive. To be a strict superset entirely containing $t_{k}$, then either $s^{\prime}<s$ and $e^{\prime}\geq e$, or $s^{\prime}\leq s$ and $e^{\prime}>e$, subject to the constraint that the width $w^{\prime}=e^{\prime}-s^{\prime}+1\leq L$. In this case, the minimum length when using the superset token $t_{k}^{\prime}$ would be:
$K_{t_{k}^{\prime}}^{s}=pl[s^{\prime}-1]+bpl[e^{\prime}+1]+1,$ (3)
which is the path length to get to the byte before $t_{k}^{\prime}$, plus the path length from the end of the data backwards to the byte after $t_{k}^{\prime}$, plus 1 for the token $t_{k}^{\prime}$ itself.
We retain a list of the widths of the tokens ending at each byte. [^6] See the expanded explanation in Appendix A for details. The set of superset tokens $S$ can be found by examining the potential $e^{\prime}$, and then seeing if the tokens ending at $e^{\prime}$ form a strict superset. Similar to the previous case, we can compute the minimum increase from replacing $t_{k}$ with a superset token by taking the minimum increase over the superset tokens $S$:
$MI_{kd}^{s}=\min_{t_{k}^{\prime}\in S}{K_{t_{k}^{\prime}}^{s}-K_{d}}.$ (4)
We then aggregate over the documents to get the overall increase for each $t\in\mathcal{V}$:
$MI_{t}=\sum_{d\in\mathcal{C}}\sum_{k=1|t_{k}=t}^{K_{d}}\min(MI_{kd}^{b},MI_{kd}^{s}).$ (5)
One iteration of this vocabulary construction procedure will have complexity $O(nL^{2})$.footnotemark:
3.3 Connecting PathPiece and Unigram
We note a connection between PathPiece and Unigram. In Unigram, the probability of a segmentation $t_{1},\dots,t_{K_{d}}$ is the product of the unigram token probabilities $p(t_{k})$:
$P(t_{1},\dots,t_{K_{d}})=\prod_{k=1}^{K_{d}}p(t_{k}).$ (6)
Taking the negative $\log$ of this product converts the objective from maximizing the likelihood to minimizing the sum of $-\log(p(t_{k}))$ terms. While Unigram is solved by the Viterbi (1967) algorithm, it can also be solved by a weighted version of PathPiece with weights of $-\log(p(t_{k}))$. Conversely, a solution minimizing the number of tokens can be found in Unigram by taking all $p(t_{k}):=1/|\mathcal{V}|$.
4 Experiments
We used the Pile corpus Gao et al. (2020); Biderman et al. (2022) for language model pre-training, which contains 825GB of English text data from 22 high-quality datasets. We constructed the tokenizer vocabularies over the MiniPile dataset Kaddour (2023), a 6GB subset of the Pile. We use the MosaicML Pretrained Transformers (MPT) decoder-only language model architecture. [^7] https://github.com/mosaicml/llm-foundry Appendix B gives the full set of model parameters, and Appendix D discusses model convergence.
4.1 Downstream Evaluation Tasks
To evaluate and analyze the performance of our tokenization process, we select 10 benchmarks from lm-evaluation-harness Gao et al. (2023). [^8] https://github.com/EleutherAI/lm-evaluation-harness These are all multiple-choice tasks with 2, 4, or 5 options, and were run with 5-shot prompting. We use arc_easy Clark et al. (2018), copa Brassard et al. (2022), hendrycksTests-marketing Hendrycks et al. (2021), hendrycksTests-sociology Hendrycks et al. (2021), mathqa Amini et al. (2019), piqa Bisk et al. (2019), qa4mre_2013 Peñas et al. (2013), race Lai et al. (2017), sciq Welbl et al. (2017), and wsc273 Levesque et al. (2012). Appendix C gives a full description of these tasks.
4.2 Tokenization Stage Variants
[…]
5 Results
Table 1 reports the downstream performance across all our experimental settings. [^10] The same table sorted by rank is in Table 9 of Appendix G. The comprehensive results for the ten downstream tasks, for each of the 350M parameter models, are given in Appendix G. A random baseline for these 10 tasks yields 32%. The Overall Avg column indicates the average results over the three vocabulary sizes. The Rank column refers to the rank of each variant with respect to the Overall Avg column (Rank 1 is best), which we will sometimes use as a succinct way to refer to a variant.
Rank Vocab Constr Init Voc Pre-tok Segment Overall Avg 32,768 Avg 40,960 Avg 49,152 Avg
1 PathPieceL BPE FirstSpace PathPieceL 49.4 49.3 49.4 49.4
9 Unigram FirstSpace 48.0 47.0 48.5 48.4
15 $n$-gram FirstSpDigit 44.8 44.6 44.9 45.0
16 $n$-gram FirstSpace 44.7 44.8 45.5 43.9
2 Unigram FirstSpace Likelihood 49.0 49.2 49.1 48.8
7 Greedy 48.3 47.9 48.5 48.6
17 PathPieceL 43.6 43.6 43.1 44.0
3 BPE FirstSpace Merge 49.0 49.0 50.0 48.1
4 Greedy 49.0 48.3 49.1 49.5
13 PathPieceL 46.5 45.6 46.7 47.2
5 WordPiece FirstSpace Greedy 48.8 48.5 49.1 48.8
6 SaGe BPE FirstSpace Greedy 48.6 48.0 49.2 48.8
8 $n$-gram FirstSpace 48.0 47.5 48.5 48.0
10 Unigram FirstSpace 47.7 48.4 46.9 47.8
11 $n$-gram FirstSpDigit 47.5 48.4 46.9 47.2
12 PathPieceR $n$-gram SpaceDigit PathPieceR 46.7 47.5 45.4 47.3
14 FirstSpDigit 45.5 45.3 45.8 45.5
18 None 43.2 43.5 44.0 42.2
Random 32.0 32.0 32.0 32.0
Table 1: Summary of 350M parameter model downstream accuracy (%). The Overall Avg column averages across the three vocabulary sizes. The Rank column refers to the Overall Avg column, best to worst.
5.1 Vocabulary Size
Figure 1: Effect of vocabulary size on downstream performance. For each tokenizer variant, we show the overall average, along with the three averages by vocabulary size, labeled according to the ranks in Table 1.
Figure 1 gives the overall average, along with the individual averages, for each of the three vocabulary sizes for each variant, labeled according to the rank from Table 1. We observe that there is a high correlation between downstream performance at different vocabulary sizes. The pairwise $R^{2}$ values for the accuracy of the 32,768 and 40,960 runs was 0.750; between 40,960 and 49,152 it was 0.801; and between 32,768 and 49,152 it was 0.834. This indicates that vocabulary size is not a crucial decision over this range of sizes. Given this high degree of correlation, we focus our analysis on the overall average accuracy. This averaging removes some of the variance amongst individual language model runs. Thus, unless specified otherwise, our analyses present performance averaged over vocabulary sizes.
5.2 Overall performance
To determine which of the differences in the overall average accuracy in Table 1 are statistically significant, we conduct a one-sided Wilcoxon signed-rank test Wilcoxon (1945) on the paired differences of the 30 accuracy scores (three vocabulary sizes over ten tasks), for each pair of variants. All $p$-values reported in this paper use this test.
Figure 2: Pairwise $p$-values for 350M model results. Black boxes lines represent $p$ > 0.05. The top 6 tokenizers are all competitive, and there is no statistically significantly best approach.
Figure 2 displays all pairwise $p$-values in a colormap. Each column designates a tokenization variant by its rank in Table 1, compared to all the ranks below it. A box is outlined in black if $p>0.05$, where we cannot reject the null. While PathPieceL-BPE had the highest overall average on these tasks, the top five tokenizers, PathPieceL-BPE, Unigram, BPE, BPE-Greedy, and WordPiece do not have any other row in Figure 2 significantly different from them. Additionally, SaGe-BPE (rank 6) is only barely worse than PathPieceL-BPE ($p$ = 0.047), and should probably be included in the list of competitive tokenizers. Thus, our first key result is that there is no tokenizer algorithm better than all others to a statistically significant degree.
All the results reported thus far are for language models with identical architectures and 350M parameters. To examine the dependency on model size, we build larger models of 1.3B parameters for six of our experiments, and 2.4B parameters for four of them. In the interest of computational time, these larger models were only run at a single vocabulary size of 40,960. The average results over the 10 task accuracies for these models are given in Figure 6 in Appendix D. See Figure 7 in Appendix D for an example checkpoint graph at each model size. The main result from these models is that the relative performance of the tokenizers does vary by model size, and that there is a group of highly-performant tokenizers that yield comparable results.
This aligns with our finding that the top six tokenizers are not statistically better than one another at the 350M model size.
5.3 Corpus Token Count vs Accuracy
Figure 3 shows the corpus token count (CTC) versus the accuracy of each vocabulary size. We do not find a straightforward relationship between the two.
Figure 3: Effect of corpus token count (CTC) vs average accuracy of individual vocabulary sizes.
The two models with the highest CTC are: PathPiece with Space pre-tokenization (12), which is to be expected given each space is its own token; and, SaGe with an initial Unigram vocabulary (10) also had a high CTC. In terms of accuracy, PathPiece with no pre-tokenization (18) and Unigram with PathPiece segmentation (17) both did quite poorly. Notably, the range of CTC is quite narrow within each vocabulary construction method, even while changes in pre-tokenization and segmentation lead to significant accuracy differences. While there are confounding factors present in this chart (e.g., pre-tokenization, vocabulary initialization, and that more tokens allow for additional computations by the downstream model) it is difficult to discern any trend that lower CTC leads to improved performance. If anything, there seems to be an inverted U-shaped curve with respect to the CTC and downstream performance.
6 Analysis
We now analyze the results across the various experiments in a more controlled manner. Our experiments allow us to examine changes in each stage of tokenization, holding the rest constant, revealing design decisions making a significant difference.
6.1 Pre-tokenization
For PathPieceR with an $n$-gram initial vocabulary, we can isolate pre-tokenization. PathPiece is efficient enough to process entire documents with no pre-tokenization, giving it full freedom to minimize the corpus token count (CTC). Adding pre-tokenization constrains PathPiece’s ability to minimize tokens, giving a natural way to vary the number of tokens. [^11] These two runs also used Digit pre-tokenization where each digit is its own token.
Figure 4 shows that PathPiece, as intended, minimizes the number of tokens used over a corpus when trained with no pre-tokenization (18). The other variants restrict spaces to either be the first character of a token (14), or their own token (12). Gallé (2019) suggests we should expect downstream performance to decrease with more tokens. Interestingly, however, we observe an increase in overall accuracy for our downstream tasks, as a function of the number of tokens. These results suggest it is not inherently beneficial to use fewer tokens, aligning with our findings in §5.3.
Figure 4: The impact of pre-tokenization on Corpus Token Count (CTC) and Overall Accuracy. Ranks in parentheses refer to performance in Table 1.
Rank Pre-tokenization Example
12 SpaceDigit The ␣ valuation ␣ is ␣ estimated ␣ to ␣ be ␣ \$ 2 1 3 M
14 FirstSpDigit The ␣valuation ␣is ␣estimated ␣to ␣be ␣\$ 2 1 3 M
18 None The ␣valu ation␣is ␣estimated ␣to␣b e␣\$ 2 1 3 M
Table 2: Example PathPiece tokenizations of “The valuation is estimated to be \$213M”; vocabulary size of 32,768.
To understand why, consider the example PathPiece tokenization in Table 2 for the three pre-tokenization methods of Figure 4. The None mode uses the word-boundary-spanning tokens “ation␣is”, “␣to␣b”, and “e␣\$”. The lack of morphological alignment demonstrated in this example is likely more important to downstream model performance than a simple token count.
Considering the $p$-values of Figure 2, SpaceDigit (12) significantly improves accuracy over FirstSpaceDigit (14) and No pre-tokenization (18). FirstSpaceDigit (14) also performed significantly better than No pre-tokenization (18).
One of our examples that isolated Digit pre-tokenization showed it was worse with significance (8 versus 11), while the other (15 versus 16) was not statistically significant. Other than mathqa, none of our downstream tasks were particularly mathematical in nature, making it difficult to conclude a definitive judgment about the utility of Digit pre-tokenization. [^12] See Figure 8 and Figure 9 in Appendix E for more details.
6.2 Vocabulary Construction
One way to examine the effects of vocabulary construction is to compare the resulting vocabularies of top-down methods trained using an initial vocabulary to the method itself. Figure 5 presents an area-proportional Venn diagram of the overlap in 40,960-sized vocabularies between BPE (6) and variants of PathPieceL (1) and SaGe (6) that were trained using an initial BPE vocabulary of size $1442^{18}=262,$ 144.131313See Figure 10 in Appendix E.2 for analogous results for Unigram, which behaves similarly. While BPE and PathPieceL overlap considerably, SaGe produces a more distinct set of tokens.
Figure 5: Venn diagram comparing 40,960 token vocabularies of BPE, PathPieceL and SaGe – the latter two were both initialized from a BPE vocabulary of 262,144.
The difference in tie breaking strategy in PathPiece between longest-token (L) and random tiebreaks (R) turns out not to be significant (see Appendix E). Further discussion of the effects of vocabulary construction through a random baseline is provided in Appendix F.
6.3 Initial Vocabulary
PathPiece, SaGe, and Unigram all require an initial vocabulary. [^14] The HuggingFace Unigram implementation starts with the one million most top $n$-grams, but sorted according to the count times the length of the token, introducing a bias toward longer tokens. For PathPiece and SaGe, we experimented with initial vocabularies of size 262,144 constructed from either the most frequent $n$-grams, or trained using either BPE or Unigram. For PathPieceL, using a BPE initial vocabulary (1) is statistically better than both Unigram (9) and $n$-grams (16), with $p\leq 0.01$. Using an $n$-gram initial vocabulary leads to the lowest performance, with statistical significance. Comparing ranks 6, 8, and 10 reveals the same pattern for SaGe, although the difference between 8 and 10 is not significant.
6.4 Segmentation
Tokenizers often use the segmentation strategy that is used in vocabulary construction. However, any vocabulary can also be used with PathPiece and with the greedy left-to-right segmentation methods.
We find that BPE works quite well with greedy segmentation (overall rank 4, insignificantly different from the top rank), but not with the shortest-path segmentation of PathPieceL (13).
Unigram, on the other hand, seems to be more tightly tied to its default maximum likelihood segmentation (2), which was significantly better than both Greedy (7) and PathPieceL (17).
7 Conclusion
We investigated the hypothesis that reducing the corpus token count (CTC) would improve downstream performance. We find that five different tokenizers, all with different CTCs, can perform comparably (§5.2). Moreover, when comparing CTC and downstream accuracy across all our experimental settings in Figure 3, we do not find a clear relationship between the two.
This casts doubt on the current understanding of why BPE is particularly effective. Our findings dovetail with Bostrom and Durrett (2020)’s results that unigram tokenizers better align with morphological segmentation compared to BPE tokenizers, further suggesting there is not a straightforward explanation of what makes a tokenizer effective.
Finally, this work contributes a number of practical contributions: (1) BPE initial vocabularies work best for top-down vocabulary constructions; (2) Space pre-tokenization Gow-Smith et al. (2022) outperforms the more common FirstSpace pre-tokenization for PathPiece; (3) overall, vocabulary size has little impact on downstream performance. To further encourage research in this direction, we make all of our tokenized vocabularies publicly available, along with the model weights from our 64 language models.
Limitations
The objective of this work is to offer a comprehensive analysis of the tokenization process. However, our findings were constrained to particular tasks and models. Given the degrees of freedom, such as choice of downstream tasks, model, vocabulary size, etc., there is a potential risk of inadvertently considering our results as universally applicable to all NLP tasks; results may not generalize to other domains of tasks.
Additionally, our experiments were exclusively with English language text, and it is not clear how these results will extend to other languages. In particular, our finding that pre-tokenization is crucial for effective downstream accuracy is not applicable to languages without space-delimited words.
We conducted experiments for three district vocabulary sizes, and we reported averaged results across these experiments. With additional compute resources and time, it could be beneficial to conduct further experiments to gain a better estimate of any potential noise. For example, in Figure 7 of Appendix D, the 100k checkpoint at the 1.3B model size is worse than expected, indicating that noise could be an issue.
Finally, the selection of downstream tasks can have a strong impact on results. To allow for meaningful results, we attempted to select tasks that were neither too difficult nor too easy for the 350M parameter models, but other choices could lead to different outcomes. There does not seem to be a good, objective criteria for selecting a finite set of task to well-represent global performance.
Ethics Statement
We have used the commonly used public dataset The Pile, which has not undergone a formal ethics review Biderman et al. (2022). Our models may include biases from the training data.
Our experimentation has used considerable energy. Each 350M parameter run took approximately 48 hours on (4) p4de nodes, each containing 8 NVIDIA A100 GPUs. We trained 62 models, including the 8 RandTrain runs in Appendix F. The (6) 1.3B parameters models took approximately 69 hours to train on (4) p4de nodes, while the (4) 2.4B models took approximately 117 hours to train on (8) p4de nodes. In total, training required 17,304 hours of p4de usage (138,432 GPU hours).
Acknowledgments
Thanks to Charles Lovering at Kensho for his insightful suggestions, and to Michael Krumdick, Mike Arov, and Brian Chen at Kensho for their help with the language model development process. This research was supported in part by the Israel Science Foundation (grant No. 1166/23).
Appendix A Expanded description of PathPiece
This section provides a self-contained explanation of PathPiece, expanding on the one in §3, with additional details on the vocabulary construction and complexity.
In order to design an optimal vocabulary $\mathcal{V}$, it is first necessary to know how the vocabulary will be used to tokenize. There can be no best vocabulary in the abstract. Thus, we first present a new lossless subword tokenizer PathPiece. This tokenization over our training corpus will provide the context to design a coherent vocabulary.
A.1 Tokenization for a given vocabulary
We work at the byte level, and require that all 256 single byte tokens are included in any given vocabulary $\mathcal{V}$. This avoids any out-of-vocabulary tokens by falling back to single bytes in the worst case.
Tokenization can be viewed as a compression problem, where we would like to tokenize text in a few tokens as possible. This has direct benefits, as it allows more text to fit in a given context window. A Minimum Description Length (MDL) argument can also be made that the tokenization using the fewest tokens best describes the data, although we saw in Subsection 6.1 this may not always hold in practice.
Tokenizers such as BPE and WordPiece make greedy decisions, such as choosing which pair of current tokens to merge to create a new one, which results in tokenizations that may use more tokens than necessary. In contrast, PathPiece will find an optimal tokenization by finding a shortest path through a Directed Acyclic Graph (DAG). Informally, each byte $i$ of training data forms a node in the graph, and there is an edge if the $w$ byte sequence ending at $i$ is a token in $\mathcal{V}$.
An implementation of PathPiece is given in Algorithm 2, where input $d$ is a text document of $n$ bytes, $\mathcal{V}$ is a given vocabulary, and $L$ is a limit on the maximum width of a token in bytes. It has complexity $O(nL)$, following directly from the two nested for-loops. It iterates over the bytes $i$ in $d$, computing 4 values for each. It computes the shortest path length $pl[i]$ in tokens up to and including byte $i$, the width $wid[i]$ of a token with that shortest path length, and the solution count $sc[i]$ of optimal solutions found thus far with that shortest length. We also remember the valid tokens of width 2 or more ending at each location $i$ in $vt[i]$, which will be used in the next section.
There will be multiple tokenizations with the same optimal length, so some sort of tiebreaker is needed. The longest token or a randomly selected token are obvious choices. We have presented the random tiebreaker method here, where a random solution is selected in a single pass in lines 29-32 of the listing using an idea from reservoir sampling Vitter (1985).
A backward pass through $d$ constructs the optimal tokenization from the $wid[e]$ values from the forward pass.
Algorithm 2 PathPiece segmentation.
1:procedure PathPiece($d,\mathcal{V},L$)
2:     $n\leftarrow\operatorname{len}(d)$triangleright document length
3:     for $i\leftarrow 1,n$ do
4:         $wid[i]\leftarrow 0$triangleright shortest path token
5:         $pl[i]\leftarrow\infty$triangleright shortest path len
6:         $sc[i]\leftarrow 0$triangleright solution count
7:         $vt[i]\leftarrow[\,]$triangleright valid token list
8:     for $e\leftarrow 1,n$ do $\triangleright$ token end
9:         for $w\leftarrow 1,L$ do $\triangleright$ token width
10:              $s\leftarrow e-w+1$triangleright token start
11:              if $s\geq 1$ then $\triangleright$ $s$ in range
12:                  $t\leftarrow d[s:e]$triangleright token
13:                  if $t\in\mathcal{V}$ then
14:                       if $s=1$ then $\triangleright$ 1 tok path
15:                           $wid[e]\leftarrow w$
16:                           $pl[e]\leftarrow 1$
17:                           $sc[e]\leftarrow 1$
18:                       else
19:                           if $w\geq 2$ then
20:                                $vt[e].\operatorname{append}(w)$
21:                           $nl\leftarrow pl[s-1]+1$
22:                           if $nl<pl[e]$ then
23:                                $pl[e]\leftarrow nl$
24:                                $wid[e]\leftarrow w$
25:                                $sc[e]\leftarrow 1$
26:                           else if $nl=pl[e]$ then
27:                                $sc[e]\leftarrow sc[e]+1$
28:                                $r\leftarrow\operatorname{rand}()$
29:                                if $r\leq 1/sc[e]$ then
30:                                    $wid[e]\leftarrow w$
31:     $T\leftarrow[\,]$triangleright output token list
32:     $e\leftarrow n$triangleright start at end of $d$
33:     while $e\geq 1$ do
34:         $w\leftarrow wid[e]$triangleright width of short path tok
35:         $s\leftarrow e-w+1$triangleright token start
36:         $t\leftarrow d[s:e]$triangleright token
37:         $T.\operatorname{append}(t)$
38:         $e\leftarrow e-w$triangleright back up a token
39:     return $\operatorname{reversed}(T)$triangleright reverse order
A.2 Optimal Vocabulary Construction
A.2.1 Vocabulary Initialization
We will build an optimal vocabulary by starting from a large initial one, and sequentially omitting batches of tokens. We start with the most frequently occurring byte $n$-grams in a training corpus, of width 1 to $L$, or a large vocabulary trained by BPE or Unigram. We then add any single byte tokens that were not already included, making room by dropping the tokens with the lowest counts. In our experiments we used an initial vocabulary size of $|\mathcal{V}|=2^{18}=262,144$.
A.2.2 Increase from omitting a token
Given a PathPiece tokenization $t_{1},\dots,t_{K_{d}}$, $\forall{d\in\mathcal{C}}$ for training corpus $\mathcal{C}$, we would like to know the increase in the overall length of a tokenization $K=\sum_{d}{K_{d}}$ from omitting a given token $t$ from our vocabulary, $\mathcal{V}\setminus\{t\}$ and recomputing the tokenization. Tokens with a low increase are good candidates to remove from the vocabulary Kudo (2018). However, doing this from scratch for each $t$ would be a very expensive $O(nL|\mathcal{V}|)$ operation.
We make a simplifying assumption that allows us to compute these increases more efficiently. We omit a specific token $t_{k}$ in the tokenization of document $d$, and compute the minimum increase $MI_{kd}$ in $K_{d}$ from not having that token $t_{k}$ in the tokenization of $d$. We then aggregate over the documents to get the overall increase for $t$:
$MI_{t}=\sum_{d\in\mathcal{C}}\sum_{k=1|t_{k}=t}^{K_{d}}MI_{kd}.$ (7)
This is similar to computing the increase from $\mathcal{V}\setminus\{t\}$, but ignores interaction effects from having several occurrences of the same token $t$ close to each other in a given document.
With PathPiece, it turns out we can compute the minimum increase in tokenization length without actually recomputing the tokenization. Any tokenization not containing $t_{k}$ must either contain a token boundary somewhere inside of $t_{k}$ breaking it in two, or it must contain a token that entirely contains $t_{k}$ as a superset. Our approach will be to enumerate all the occurrences for these two cases, and to find the minimum increase $MI_{kd}$ overall.
Before considering these two cases, there is a shortcut that often tells us that there would be no increase due to omitting $t_{k}$ ending at index $e$. We computed the solution count vector $sc[e]$ when running Algorithm 2. If $sc[e]>1$ for a token ending at $e$, then the backward pass could simply select one of the alternate optimal tokens, and find an overall tokenization of the same length.
Let $t_{k}$ start at index $s$ and end at index $e$, inclusive. Remember that path length $pl[i]$ represents the number of tokens required for shortest path up to and including byte $i$. We can also run Algorithm 2 backwards on $d$, computing a similar vector of backwards path lengths $bpl[i]$, representing the number of tokens on a path from the end of the data up to and including byte $i$. The overall minimum length of a tokenization with a token boundary after byte $j$ is thus:
$K_{j}^{b}=pl[j]+bpl[j+1].$ (8)
We have added an extra constraint on the shortest path, that there is a break at $j$, so clearly $K_{j}^{br}\geq pl[n]$. The minimum increase for the case of having a token boundary within $t_{k}$ is thus:
$MI_{kd}^{b}=\min_{j=s,\dots,e-1}{K_{j}^{b}-pl[n]}.$ (9)
Each token $t_{k}$ will have no more than $L-1$ potential internal breaks, so the complexity of computing $MI_{kd}^{b}$ is $O(L)$.
The minimum increase from omitting $t_{k}$ could also be on a tokenization containing a strict superset of $t_{k}$. Let this superset token be $t_{k}^{\prime}$, with start $s^{\prime}$ and end $e^{\prime}$ inclusive. To be a strict superset jumping over $t_{k}$, we must have $s^{\prime}<s$ and $e^{\prime}\geq e$, or $s^{\prime}\leq s$ and $e^{\prime}>e$, subject to the constraint that the width $w^{\prime}=e^{\prime}-s^{\prime}+1\leq L$. In this case, the minimum length of using the superset token $t_{k}^{\prime}$ would be:
$K_{t_{k}^{\prime}}^{s}=pl[s^{\prime}-1]+bpl[e^{\prime}+1]+1,$ (10)
which is the path length to get to the byte before $t_{k}^{\prime}$, plus the path length go backwards to the byte after $t_{k}^{\prime}$, plus 1 for the token $t_{k}^{\prime}$ itself.
We remembered a list of the widths of the tokens ending at each byte, $vt[e]$ in Algorithm 2. The set of superset tokens $S$ can be found by examining the $O(L)$ potential $e^{\prime}$, and then seeing if the $w^{\prime}\in vt[e^{\prime}]$ give tokens forming a strict superset. There are $O(L)$ potential tokens ending at $e^{\prime}$ in $vt[e^{\prime}]$, so the overall complexity of finding the superset tokens is $O(L^{2})$
Similar to the previous case, we can compute the minimum increase from replacing $t_{k}$ with a superset token by taking the minimum increase over the superset tokens:
$MI_{kd}^{s}=\min_{t_{k}^{\prime}\in S}{K_{t_{k}^{\prime}}^{s}-pl[n]}.$ (11)
Finally, the overall minimum increase $MI_{kd}$ from omitting $t_{k}$ is simply
$MI_{kd}=\min(MI_{kd}^{b},MI_{kd}^{s}).$ (12)
When aggregating over all $t_{k}$ according to eq (7), one iteration of the vocabulary construction procedure will have complexity $O(nL^{2})$.
Appendix B Language Model Parameters
The 350M parameter models were trained using the MPT architecture [^15] https://github.com/mosaicml/llm-foundry with the following parameters:
# Modelmodel:  name: mpt_causal_lm  init_deice: meta  d_model: 1024  n_heads: 16  n_layers: 24  expansion_ratio: 4  max_seq_len: 2048  attn_config:    alibi: true    attn_impl: triton    clip_qkv: 6# Optimizationdevice_eval_batch_size: 5device_train_microbatch_size: 32global_train_batch_size: 1024 # ~2M tokensmax_duration: 100000ba # ~200B tokensoptimizer:  name: decoupled_adamw  lr: 3.0e-4  betas:  - 0.9  - 0.95  eps: 1.0e-08  weight_decay: 0.0001scheduler:  name: cosine_with_warmup  t_warmup: 0.05dur  alpha_f: 0.1# Systemprecision: amp_bf16# Algos and Callbacksalgorithms:  gradient_clipping:    clipping_threshold: 1    clipping_type: norm
The 1.3B parameter models simply changes:
d_model: 1024
The 2.4B parameter models updates:
d_model: 2560n_heads: 20n_layers: 32
Appendix C Description of Downstream Tasks
To evaluate the performance of our various tokenization experiments, we select ten competitive benchmarks from lm-evaluation-harness Gao et al. (2023) [^16] https://github.com/EleutherAI/lm-evaluation-harness, that we broadly categorize into three types of Question Answering (QA) tasks: Knowledge-based, Common-sense Reasoning and Context-based.
Knowledge Based Tasks Knowledge based tasks in this study expect LLMs to answer questions based on domain-specific internal retrieval. Our Knowledge-based baselines in this work include:
SciQ: The SciQ task, proposed by Welbl et al. (2017) contains a total of 13,679 science exam questions. The questions are in multiple-choice format with 4 answer options each. An additional text is provided as supporting evidence for a majority of the answers.
ARC (AI2 Reasoning Challenge): Clark et al. (2018) compiles grade-school level, multiple-choice science question dataset consists of 7,787 science exam questions that are split into “easy” and “hard” sets. For this study, we employ the easy set of 5,197 questions, each having 4 answer choices.
MathQA: Amini et al. (2019) introduce a dataset of math word problems that require LLMs to use their internal understanding of mathematical equations and arithmetic comprehension. Similar to SciQ, this dataset consists of 37k multiple-choice questions with the equations for each used annotated.
HendrycksTest: Hendrycks et al. (2021) provide a comprehensive suite of of multiple-choice tests for assessing text models in multi-task contexts. It comprises of 57 tasks such as elementary mathematics, US history, law of which we use the sociology and marketing tests.
Commonsense Reasoning Tasks These tasks assess the model’s capability to infer and reason about everyday scenarios based on implicit knowledge.
COPA (Choice of Plausible Alternatives): COPA proposed by Brassard et al. (2022) is a benchmark for assessing progress in open-domain commonsense causal reasoning. It consists of 1000 questions where each question is composed of a premise and two alternatives. The task is to select the alternative that more plausibly has a causal relation with the premise.
PiQA (Physical Interaction Question Answering): Bisk et al. (2019) introduce a task that assess the understanding of physical commonsense by language models. Comprised of everyday situation with a preference for atypical solutions, this dataset is formulated as multiple choice question with two possible solutions choices for each question.
Winograd Schema Challenge: Levesque et al. (2012) define a task with a pair of sentences that differ only in one or two words and that contain a referential ambiguity that is resolved in opposite directions in the two sentences. This dataset of 273 tasks test language model understanding of the content of the text and disambiguation ability.
Context Based Tasks These tasks are reliant on understanding context and drawing conclusions from it.
RACE (Reading Comprehension from Examinations): RACE proposed by Lai et al. (2017) is a collection of English questions set aside to Chinese school students. Each item is divided into two parts, a passage that the student must read and a set of 4 potential answers, requiring extraction and reasoning capabilities.
QA4MRE (Question Answering for Machine Reading Evaluation): QA4MRE by Peñas et al. (2013) is a benchmark designed to resolve reading comprehension challenges. This task focuses on reading of single documents and identifying the answers to a set of questions. Questions are in the form of multiple choice with one correct option.
Our goal was to select tasks where a 350M parameter model could do significantly better than random chance, avoiding evaluation right at the noisier random threshold. We started with the tasks that had a non-zero random score (indicating multiple choice), and then chose tasks where BPE at a vocabulary size 40,960 could do well above random. In the end, the average accuracy across models was more than 15% above random on all tasks.
Note that in results tables we have shortened the name hendrycksTest-marketing to marketing, hendrycksTest-sociology to sociology, and qa4mre_2013 to qa4mre.
Appendix D Effect of model size and model convergence
To examine the dependency on model size, we build larger models of 1.3B parameters for 6 of our experiments, and 2.4B parameters for 4 of them. These models were trained over the same 200 billion tokens. In the interest of computational time, these larger models were only run at a single vocabulary size of 40,960. The average results over the 10 task accuracies for these models is given in Figure 6. See Table 12 in Appendix G for the numerical values.
Figure 6: 40,960 vocab average accuracy at various models sizes
It is noteworthy from the prevalence of crossing lines in Figure 6 that the relative performance of the tokenizers do vary by model size, and that there is a group of tokenizers that are trading places being at the top for various model sizes. This aligns with our observation that the top 6 tokenizers were within the noise, and not significantly better than each other in the 350M models.
Each model was trained on around 200 billion tokens. Figure 7 gives a plot of the average accuracy for PathPieceL with a BPE initial vocabulary and a vocabulary size of 40,960 at various checkpoints in the language model training process. It also shows checkpoints for the larger 1.3B and 2.4B models discussed in the Limitations section. With the exception of the 100k checkpoint at 1.3B, the model appears to be continually improving. It is unclear why the 100k checkpoint did so poorly.
Figure 7: Checkpoint accuracy values for PathPieceL with an initial vocabulary from BPE and a vocabulary size of 40,960, evaluated at 5 checkpoints.
Appendix E Additional Analysis
Here we additional details for results from §6 that are just summarized in the text in the interest of space.
E.1 Digit Pre-tokenization
We have two examples isolating Digit pre-tokenization, when a digit must always be its own token. Figure 8 shows Digit hurts for Sage with an $n$-gram initial vocabulary, while Figure 9 shows no significant differences for PathPieceL, also with an $n$-gram initial vocabulary.
Figure 8: Pre-tokenization of Sage, $n$-gram initial,
$p$=0.025.
Figure 9: Pre-tokenization of PathPieceL $n$-gram,
$p$=0.54.
With the exception of mathqa, none of our downstream tasks were particularly mathematical in nature. It is likely this makes it hard to make a definitive judgement on Digit with our experiments.
E.2 Vocabulary Construction
Figure 10 gives a Venn diagram of the overlap in vocabularies between Unigram, PathPieceL, and SaGe, when both PathPieceL and SaGe were constructed from a large initial vocabulary of size 262,144 from Unigram. As with Figure 5, we see that PathPiece is more similar to Unigram, while SaGe chose more distinct tokens.
Figure 10: Venn diagrams comparing 40,960 token vocabularies of Unigram, PathPieceL and SaGe, where the latter two were both trained from a initial Unigram vocabulary of size 262,144
E.3 PathPiece tie breaking
The difference in tie breaking between choosing the longest token with PathPieceL versus choosing randomly with PathPieceR turns out not to be significant, as seen in in Figure 11.
Figure 11: Tiebreaking PathPieceL vs PathPieceR
with $n$-gram, $p$=0.067.
Appendix F RandTrain
None of our experiments completely isolate the effect of the vocabulary construction step. We created a new baseline random vocabulary construction approach, RandTrain, in an attempt to do so. It is meant to work with a top-down method like SaGe or PathPieceL, and uses the same initial vocabulary, pre-tokenization, and segmentation as either of those, with a simple vocabulary construction algorithm.
We compute a count for each token in the vocabulary. For the top $n$-gram initial vocabulary it is simply the $n$-gram count from the training corpus. For a BPE initial vocabulary we tokenized the training corpus with BPE and the large initial vocabulary, and then use the occurrence counts of each token. We normalize these counts into target selection probabilities $p_{k}$ for token $t_{k}$.
The RandTrain vocabulary construction process is simply to randomly sample our desired vocabulary size $m$ of tokens from the initial vocabulary, proportionally to $p_{k}$, without replacement. Sampling without replacement is necessary to avoid have duplicate words in the vocabulary. Interestingly, this is not possible if there are any $p_{k}>1/m$, which are termed infeasible or overweight items Efraimidis (2010). The intuition behind this is when selecting $m$ items without replacement, it is not possible to select a given item more than once. So even if an item is always selected in a sample, the selection probability will be $p_{k}=1/m$.
We sampled without replacement using the A-ES Algorithm described in Efraimidis (2010). A significant number the most common tokens in the vocabulary were infeasible and hence were unable to reach their target $p_{k}$. A token with a higher $p_{k}$ is more likely to be sampled than a token with a lower one, but they may significantly differ from their target $p_{k}$.
We build 6 RandTrain models with 3 different types of pre-tokenization, and with Greedy segmentation to compare to SaGe, and PathPieceL segmentation to compare to PathPieceL. We only used a single vocabulary size of 40,960, so $p$-values are only computed on the 10 task accuracies, rather than the 30 used elsewhere. Task level accuracies are given in Table 5 and Table 6 in Appendix G.
Before comparing RandTrain to SaGe and PathPieceL, we will compare our RandTrain runs to each other, with different segmentation approaches. In Figure 12 and Figure 14 we have pairs of RandTrain runs that only vary by the segmentation method.
Figure 12: Comparison of Greedy and PathPieceL segmentation, with RandTrain vocabulary construction, BPE initial vocab, and FirstSpace pre-tokenization, $p$=0.0273
Figure 13: Comparison of Greedy and PathPieceL segmentation, with RandTrain vocabulary construction, $n$-gram initial vocab, and FirstSpace pre-tokenization, $p$=0.00195
Figure 14: Comparison of Greedy and PathPieceL segmentation, with RandTrain vocabulary construction, $n$-gram initial vocab, and FirstSpaceDigit pre-tokenization, $p$=0.00293
In line with Subsection 6.4, Greedy performs significantly better than PathPieceL segmentation in all 3 cases. However, for the two cases with an $n$-gram initial vocabulary the PathPieceL segmentation did extremely poorly. The RandTrain vocabulary construction, $n$-gram initial vocabulary, and PathPieceL segmentation interact somehow to give accuracies well below any others.
This makes the comparison of RandTrain to PathPieceL less informative. We can see in Figure 15 that PathPieceL is significantly better than RandTrain with a BPE initial vocabulary.
However, the other two comparisons in Figure 16 are Figure 17 are not that meaningful. They are significantly better, but that is more about the weak baseline of RandTrain with PathPieceL segmentation than anything positive about PathPieceL.
Figure 15: Comparison of PathPieceL and RandTrain, with BPE initial vocab, and FirstSpace pre-tokenization, $p$=0.0137
Figure 16: Comparison of PathPieceL and RandTrain, with $n$-gram initial vocab, and FirstSpace pre-tokenization, $p$=9.77e-4
Figure 17: Comparison of PathPieceL and RandTrain, with $n$-gram initial vocab, and FirstSpaceDigits pre-tokenization, $p$=0.00977
The remaining comparison between SaGe and RandTrain is more interesting. In Figure 18 and Figure 19 SaGe was not significantly better than RandTrain, with a $p$-value of 0.0645.
Figure 18: Comparison of SaGe and RandTrain, with BPE initial vocab, and FirstSpace pre-tokenization, $p$=0.0645
The cases is even worse for the two $n$-gram initial vocabulary cases. In Figure 19 the $p$-value was a 0.688, and in Figure 20 RandTrain was actually better, although not significantly.
Figure 19: Comparison of SaGe and RandTrain, with $n$-gram initial vocab, and FirstSpace pre-tokenization, $p$=0.688
Figure 20: Comparison of RandTrain and SaGe, with $n$-gram initial vocab, and FirstSpaceDigit pre-tokenization, $p$=0.15
We saw in Table 1 that both PathPieceL-BPE and SaGe-BPE are effective tokenizers. In attempting to isolate the benefit from the vocabulary construction step, we see that PathPieceL-BPE outperforms our simple baseline. However, SaGe was unable to outperform the baseline, perhaps implying that RandTrain may actually be a simple but fairly effective vocabulary construction method.
Appendix G Detailed Experimental Results
This section gives the detailed accuracy results for the 10 downstream evaluation tasks on each model that was trained. The tables are divided by the vocabulary size used, with Table 3 and Table 4 for 32,768; Table 5 and Table 6 for 40,960; and Table 7 and Table 8 for 49,152. The highest value or values (in the case of ties) are shown in bold. Table 9 show the same results as Table 1, but are sorted from best to worst by rank.
The detailed accuracy results for our 1.3B parameter models, which were all performed at a single vocabulary size of 40,960, are given in Table 10 and Table 11. Average accuracy results for larger models of 1.3B and 2.4B parameters are given in Table 12. See §Limitations for more discussion of this table.
Vocab Constr Init Voc Pre-tok Segmentation average arc_easy copa marketing mathqa piqa
BPE FirstSpace Merge 48.8 51.2 69.0 32.9 23.9 66.3
FirstSpace Greedy 48.3 51.9 66.0 32.9 23.7 65.6
FirstSpace PathPieceL 45.6 45.6 61.0 29.9 23.0 60.5
Unigram FirstSpace Likelihood 49.2 50.7 73.0 30.8 23.1 66.3
FirstSpace Greedy 47.9 50.3 68.0 31.2 23.1 65.2
FirstSpace PathPieceL 43.6 41.2 57.0 31.6 22.0 60.6
WordPiece FirstSpace Greedy 48.5 52.5 64.0 32.5 23.9 65.6
SaGe BPE FirstSpace Greedy 47.9 49.7 67.0 26.5 23.2 65.9
$n$-gram FirstSpDigit Greedy 48.4 50.3 71.0 29.5 22.0 65.1
$n$-gram FirstSpace Greedy 47.5 48.8 64.0 29.5 23.0 66.6
Unigram FirstSpace Greedy 48.4 52.0 74.0 27.8 22.7 65.7
PathPieceL BPE FirstSpace PathPieceL 49.3 50.8 68.0 34.2 23.0 66.4
$n$-gram FirstSpace PathPieceL 44.8 42.3 61.0 27.4 23.0 61.2
$n$-gram FirstSpDigit PathPieceL 44.6 42.3 62.0 31.2 22.8 61.2
Unigram FirstSpace PathPieceL 46.9 50.4 64.0 24.8 23.5 66.2
PathPieceR $n$-gram FirstSpDigit PathPieceR 45.3 46.9 67.0 26.9 22.4 59.9
$n$-gram None PathPieceR 43.5 42.5 65.0 26.1 22.8 61.7
$n$-gram SpaceDigit PathPieceR 47.5 48.6 68.0 32.9 23.3 65.0
Random 32.0 25.0 50.0 25.0 20.0 50.00
Table 3: 350M parameter model, 32,768 token vocabulary, accuracy (%) on average and initial 5 tasks
Vocab Constr Init Voc Pre-tok Segmentation qa4mre race sciq sociology wsc273
BPE FirstSpace Merge 29.6 29.2 87.3 30.9 67.8
FirstSpace Greedy 27.5 30.7 88.0 30.9 66.3
FirstSpace PathPieceL 28.2 29.0 83.8 28.4 66.3
Unigram FirstSpace Likelihood 31.0 30.2 86.4 31.8 68.5
FirstSpace Greedy 28.9 30.6 86.9 31.8 62.6
FirstSpace PathPieceL 29.9 27.5 74.6 26.4 65.6
WordPiece FirstSpace Greedy 32.0 30.7 88.5 27.9 67.4
SaGe BPE FirstSpace Greedy 31.7 30.2 89.0 28.4 67.8
$n$-gram FirstSpDigit Greedy 31.0 30.3 86.6 32.3 66.0
$n$-gram FirstSpace Greedy 30.0 31.0 87.8 25.9 68.5
Unigram FirstSpace Greedy 29.6 28.9 88.2 32.3 63.0
PathPieceL BPE FirstSpace PathPieceL 28.5 31.1 88.8 35.3 67.0
$n$-gram FirstSpace PathPieceL 30.3 27.3 80.0 32.8 62.6
$n$-gram FirstSpDigit PathPieceL 27.8 25.5 79.2 31.3 62.6
Unigram FirstSpace PathPieceL 29.6 30.6 87.6 24.4 68.1
PathPieceR $n$-gram FirstSpDigit PathPieceR 28.5 29.4 78.6 28.9 64.5
$n$-gram None PathPieceR 27.1 27.0 77.7 28.9 56.0
$n$-gram SpaceDigit PathPieceR 25.0 29.4 85.7 32.3 64.8
Random 25.0 25.0 25.0 25.0 50.0
Table 4: 350M parameter model, 32,768 token vocabulary, accuracy (%) on remaining 5 tasks
Vocab Constr Init Voc Pre-tok Segmentation average arc_easy copa marketing mathqa piqa
BPE FirstSpace Merge 50.0 52.7 70.0 31.6 24.3 66.9
FirstSpace Greedy 49.1 52.3 66.0 27.4 22.9 66.9
FirstSpace PathPieceL 46.7 48.0 58.0 27.4 23.4 62.1
Unigram FirstSpace Likelihood 49.1 51.4 71.0 32.1 23.4 66.1
Unigram FirstSpace Greedy 48.5 49.9 64.0 30.3 23.3 65.7
Unigram FirstSpace PathPieceL 43.1 40.5 56.0 28.6 23.0 60.3
WordPiece FirstSpace Greedy 49.1 52.3 70.0 28.6 23.7 66.5
SaGe BPE FirstSpace Greedy 49.2 50.8 70.0 29.9 23.2 66.4
$n$-gram FirstSpDigit Greedy 46.9 48.4 67.0 30.3 22.6 64.0
$n$-gram FirstSpace Greedy 48.5 49.8 68.0 32.9 22.8 65.4
Unigram FirstSpace Greedy 46.9 51.7 65.0 28.6 23.9 65.2
PathPieceL BPE FirstSpace PathPieceL 49.4 52.1 71.0 29.9 23.9 66.9
$n$-gram FirstSpace PathPieceL 45.5 42.6 63.0 30.3 22.7 60.9
$n$-gram FirstSpDigit PathPieceL 44.9 44.0 60.0 29.9 22.6 60.8
Unigram FirstSpace PathPieceL 48.5 51.7 71.0 31.2 24.2 66.2
PathPieceR $n$-gram FirstSpDigit PathPieceR 45.8 47.5 63.0 28.2 22.4 60.7
$n$-gram None PathPieceR 44.0 41.2 66.0 26.5 21.6 62.4
$n$-gram SpaceDigit PathPieceR 45.4 46.3 64.0 32.1 22.7 60.0
RandTrain BPE FirstSpace Greedy 48.6 50.5 70.0 29.5 23.4 65.8
$n$-gram FirstSpDigit Greedy 47.9 50.0 63.0 29.5 23.3 65.3
$n$-gram FirstSpace Greedy 48.3 50.3 70.0 28.2 24.3 65.8
$n$-gram None Greedy 42.2 41.3 55.0 27.4 21.7 63.2
BPE FirstSpace PathPieceL 46.5 45.8 65.0 30.8 23.3 62.8
$n$-gram FirstSpDigit PathPieceL 38.8 31.2 48.0 27.8 22.6 54.7
$n$-gram FirstSpace PathPieceL 40.0 30.7 55.0 26.5 20.8 55.4
$n$-gram None PathPieceL 36.8 27.7 56.0 28.6 22.8 54.5
random 32.0 25.0 50.0 25.0 20.0 50.0
Table 5: 350M parameter model, 40,960 token vocabulary, accuracy (%) on average and initial 5 tasks
Vocab Constr Init Voc Pre-tok Segmentation qa4mre race sciq sociology wsc273
BPE FirstSpace Merge 32.4 30.1 87.7 35.3 69.2
FirstSpace Greedy 31.7 30.9 88.3 35.8 68.9
FirstSpace PathPieceL 30.3 30.2 83.8 35.3 68.1
Unigram FirstSpace Likelihood 29.6 30.8 86.4 32.8 67.8
FirstSpace Greedy 32.4 29.6 86.7 32.8 70.3
FirstSpace PathPieceL 30.3 27.4 75.0 27.4 62.3
WordPiece FirstSpace Greedy 31.0 30.3 87.7 32.8 68.1
SaGe BPE FirstSpace Greedy 28.9 30.2 89.5 34.8 67.8
$n$-gram FirstSpDigit Greedy 30.6 28.1 85.8 32.3 59.7
$n$-gram FirstSpace Greedy 29.2 30.0 88.4 33.3 65.2
Unigram FirstSpace Greedy 26.8 29.1 86.9 31.3 60.1
PathPieceL BPE FirstSpace PathPieceL 31.0 29.6 87.3 34.3 67.8
$n$-gram FirstSpace PathPieceL 29.9 27.9 81.0 34.8 61.9
$n$-gram FirstSpDigit PathPieceL 27.5 28.2 80.7 30.9 64.1
Unigram FirstSpace PathPieceL 31.3 29.7 86.3 29.9 63.7
PathPieceR $n$-gram FirstSpDigit PathPieceR 29.9 30.8 82.1 27.4 66.3
$n$-gram None PathPieceR 23.6 28.3 73.8 35.8 60.4
$n$-gram SpaceDigit PathPieceR 27.5 28.7 78.2 31.3 63.0
RandTrain BPE FirstSpace Greedy 32.0 29.6 86.9 30.9 67.4
$n$-gram FirstSpDigit Greedy 30.6 30.0 87.5 31.3 68.1
$n$-gram FirstSpace Greedy 29.9 29.7 85.3 32.8 67.0
$n$-gram None Greedy 28.2 27.8 75.9 26.4 55.0
BPE FirstSpace PathPieceL 32.8 28.5 80.3 30.9 64.5
$n$-gram FirstSpDigit PathPieceL 31.3 24.2 62.1 30.4 55.3
$n$-gram FirstSpace PathPieceL 28.9 23.6 66.8 33.8 59.0
$n$-gram None PathPieceL 21.5 24.9 51.8 28.9 51.7
random 25.0 25.0 25.0 25.0 50.0
Table 6: 350M parameter model, 40,960 token vocabulary, accuracy (%) on remaining 5 tasks
Vocab Constr Init Voc Pre-tok Segmentation average arc_easy copa marketing mathqa piqa
BPE FirstSpace Merge 48.1 52.3 65.0 31.6 23.7 65.7
FirstSpace Greedy 49.5 53.9 72.0 31.6 24.2 68.4
FirstSpace PathPieceL 47.2 48.6 69.0 26.9 22.8 63.1
Unigram FirstSpace Likelihood 48.8 52.3 69.0 35.0 23.9 66.1
FirstSpace Greedy 48.6 51.6 68.0 32.1 24.4 65.7
FirstSpace PathPieceL 44.0 39.4 57.0 30.3 23.3 61.2
WordPiece FirstSpace Greedy 48.8 52.6 68.0 28.2 23.5 66.2
SaGe BPE FirstSpace Greedy 48.8 51.9 71.0 29.9 22.6 65.5
$n$-gram FirstSpDigit Greedy 47.2 46.6 67.0 31.2 22.7 63.4
$n$-gram FirstSpace Greedy 48.0 49.7 66.0 31.6 21.6 65.7
Unigram FirstSpace Greedy 47.8 49.7 68.0 29.9 23.5 64.6
PathPieceL BPE FirstSpace PathPieceL 49.4 51.9 69.0 29.9 24.5 66.6
$n$-gram FirstSpace PathPieceL 43.9 42.4 56.0 28.6 23.8 60.3
$n$-gram FirstSpDigit PathPieceL 45.0 44.5 59.0 28.2 22.3 59.5
Unigram FirstSpace PathPieceL 48.4 51.4 67.0 29.5 24.7 65.2
PathPieceR $n$-gram FirstSpDigit PathPieceR 45.5 46.0 62.0 25.6 22.1 61.6
$n$-gram None PathPieceR 42.2 42.6 64.0 22.2 22.4 60.9
$n$-gram SpaceDigit PathPieceR 47.3 48.7 68.0 34.2 21.9 65.1
random 32.0 25.0 50.0 25.0 20.0 50.0
Table 7: 350M parameter model, 49,152 token vocabulary, accuracy (%) on average and initial 5 tasks
Vocab Constr Init Voc Pre-tok Segmentation qa4mre race sciq sociology wsc273
BPE FirstSpace Merge 28.9 31.0 87.3 28.9 67.0
FirstSpace Greedy 29.6 31.2 88.4 29.4 66.3
FirstSpace PathPieceL 31.0 30.7 85.4 31.8 63.0
Unigram FirstSpace Likelihood 27.5 30.3 89.1 28.9 65.9
FirstSpace Greedy 32.4 29.5 86.7 32.3 63.7
FirstSpace PathPieceL 33.1 26.0 74.5 27.9 67.0
WordPiece FirstSpace Greedy 29.2 31.1 88.0 34.3 66.7
SaGe BPE FirstSpace Greedy 29.6 31.2 87.5 32.3 65.9
$n$-gram FirstSpDigit Greedy 29.2 28.8 86.4 34.3 61.9
$n$-gram FirstSpace Greedy 28.8 30.2 87.5 33.8 64.5
Unigram FirstSpace Greedy 28.9 31.4 87.0 29.9 65.6
PathPieceL BPE FirstSpace PathPieceL 31.0 31.4 87.5 31.3 70.7
$n$-gram FirstSpace PathPieceL 27.5 26.7 80.8 32.3 60.8
$n$-gram FirstSpDigit PathPieceL 28.9 30.0 80.6 35.8 61.2
Unigram FirstSpace PathPieceL 29.2 30.5 88.5 32.8 65.6
PathPieceR $n$-gram FirstSpDigit PathPieceR 29.6 29.5 82.8 30.9 64.5
$n$-gram None PathPieceR 25.7 27.5 72.5 27.4 57.1
$n$-gram SpaceDigit PathPieceR 27.5 28.7 84.0 28.9 66.3
Random 25.0 25.0 25.0 25.0 50.0
Table 8: 350M parameter model, 49,152 token vocabulary, accuracy (%) on remaining 5 tasks
Rank Vocab Constr Init Voc Pre-tok Segment Overall avg 32,768 avg 40,960 avg 49,152 avg
1 PathPieceL BPE FirstSpace PathPieceL 49.4 49.3 49.4 49.4
2 Unigram FirstSpace Likelihood 49.0 49.2 49.1 48.8
3 BPE FirstSpace Merge 49.0 48.8 50.0 48.1
4 BPE FirstSpace Greedy 49.0 48.3 49.1 49.5
5 WordPiece FirstSpace Greedy 48.8 48.5 49.1 48.8
6 SaGe BPE FirstSpace Greedy 48.6 47.9 49.2 48.8
7 Unigram FirstSpace Greedy 48.3 47.9 48.5 48.6
8 SaGe $n$-gram FirstSpace Greedy 48.0 47.5 48.5 48.0
9 PathPieceL Unigram FirstSpace PathPieceL 48.0 46.9 48.5 48.4
10 SaGe Unigram FirstSpace Greedy 47.7 48.4 46.9 47.8
11 SaGe $n$-gram FirstSpDigit Greedy 47.5 48.4 46.9 47.2
12 PathPieceR $n$-gram SpaceDigit PathPieceR 46.7 47.5 45.4 47.3
13 BPE FirstSpace PathPieceL 46.5 45.6 46.7 47.2
14 PathPieceR $n$-gram FirstSpDigit PathPieceR 45.5 45.3 45.8 45.5
15 PathPieceL $n$-gram FirstSpDigit PathPieceL 44.8 44.6 44.9 45.0
16 PathPieceL $n$-gram FirstSpace PathPieceL 44.7 44.8 45.5 43.9
17 Unigram FirstSpace PathPieceL 43.6 43.6 43.1 44.0
18 PathPieceR $n$-gram None PathPieceR 43.2 43.5 44.0 42.2
Random 32.0 32.0 32.0 32.0
Table 9: Summary of 350M parameter model downstream accuracy (%), sorted by rank
Vocab Constr Init Voc Pre-tok Segmentation average arc_easy copa marketing mathqa piqa
BPE FirstSpace Merge 53.1 62.0 77.0 32.1 25.0 71.1
Unigram FirstSpace Likelihood 52.4 60.6 71.0 30.3 25.2 71.0
SaGe BPE FirstSpace Greedy 52.2 62.0 72.0 27.4 24.5 71.6
$n$-gram FirstSpDigit Greedy 50.7 60.3 71.0 28.6 22.8 69.4
PathPieceL BPE FirstSpace PathPieceL 49.2 57.4 66.0 27.8 24.3 65.9
$n$-gram FirstSpDigit PathPieceL 47.6 49.7 67.0 24.8 23.4 63.2
$n$-gram SpaceDigit PathPieceL 46.3 51.1 59.0 28.6 23.3 63.8
Random 32.0 25.0 50.0 25.0 20.0 50.0
Table 10: 1.3B parameter model, 40,960 token vocabulary, accuracy (%) on average and initial 5 tasks
Vocab Constr Init Voc Pre-tok Segmentation qa4mre race sciq sociology wsc273
BPE FirstSpace Merge 32.4 34.9 93.0 26.4 76.9
Unigram FirstSpace Likelihood 37.7 33.0 91.8 28.9 74.4
SaGe BPE FirstSpace Greedy 34.9 34.8 92.5 25.9 76.2
$n$-gram FirstSpDigit Greedy 29.9 32.9 91.5 29.4 71.1
PathPieceL BPE FirstSpace PathPieceL 31.0 33.3 89.4 26.4 70.7
$n$-gram FirstSpDigit PathPieceL 31.0 31.6 86.1 29.4 70.0
$n$-gram SpaceDigit PathPieceL 28.9 31.3 87.1 22.4 67.0
Random 25.0 25.0 25.0 25.0 50.0
Table 11: 1.3B parameter model, 40,960 token vocabulary, accuracy (%) on remaining 5 tasks
Voc Con Init V Pre-tok Seg 350M avg 350M rnk 1.3B avg 1.3B rnk 2.4B avg 2.4B rnk
BPE FirSp Merge 50.0 1 53.1 1 54.2 3
PathPL BPE FirSp PathPL 49.4 3 49.2 5 52.7 4
PathPL $n$-gram FirSpD PathPL 44.9 6 47.6 6
SaGe BPE FirSp Greedy 49.2 2 52.2 3 55.0 1
SaGe $n$-gram FirSpD Greedy 46.9 5 50.7 4
Unigram FirSp Likeli 49.1 4 52.4 2 54.7 2
Table 12: Downstream accuracy (%) of 10 tasks with vocab size 40,960, for various model sizes